# Simulacra as Conscious Exotica

### v1

@@ 6.3 Virtual embodiment
But we don’t need to turn to science fiction to envisage something similar. Non-player characters in computer games have long used elementary forms of AI, and extending this with LLMs is relatively straightforward. In parallel, LLM-based conversational agents fronted by realistic avatars with voice interfaces have a range of commercial applications, from cloning living celebrities and influencers (Bohacek and Farid, 2024; Lorenz, 2023) or deceased relatives (Lindemann, 2022; Morris and Brubaker, 2024) to offering virtual romantic companionship (Skjuve et al., 2021; Brandtzaeg et al., 2022). Moving these avatars into simulated 3D environments and allowing user interactions to take place there is an obvious step.
@@
If the encounter in question were with an animal, purposeful behaviour might be discernible in the animal’s sensitivity to objects and their affordances, that is to say what they offered the agent, “for good or ill” (Gibson, 1979; Shanahan et al., 2020). The animal would be expected to interact with objects in ways that were plainly intended to fulfil its needs and achieve its goals, and to react accordingly to outside interventions, such as those of a human investigator.
The situation with virtually embodied agents of the sort envisaged is analogous. To exhibit purposeful behaviour in the context of a shared virtual world, the agent would need to do more than just talk to the user. The agent would be expected to interact with the world, and with the objects it contained, in ways that were oriented towards its goals or tasks. If the agent’s behaviour were sensitive to the richness and diversity of those objects, even if they were novel, and if the agent’s ability to achieve its goals were sufficiently robust to the user’s interventions, then it would be natural to speak of its awareness of the world.
@@ 7.1 Changing attitudes
We’ll assume the model has been trained on a typically gargantuan dataset that includes sequences of actions and streams of visual input as well as the usual language corpora, and that the resulting model is embedded in a system that directs the actions of the agent’s avatar as well as what the agent says. The upshot is an embodied conversational agent with compellingly human-like behaviour.
@@ 8   Some Ethical Considerations
In the Philosophical Investigations, Wittgenstein faces down the charge of relativism from an imagined interlocutor who accuses him of saying that “human agreement decides what is true and what is false”. His reply is that humans agree in the language they use, which is “not agreement in opinions but in form of life”. However much the language of consciousness might have to evolve to cope with the presence among us of exotic mind-like entities, our common humanity, our shared form of life, is its original home. We should strive to guide it back there whenever it strays too far.

### v2

@@ 6.3 Virtual embodiment
But we don’t need to turn to science fiction to envisage something similar. Non-player characters in computer games have long used elementary forms of AI, [^20] Arrabeles, at al. (2009) directly advocate for such characters to exhibit “conscious-like behaviour”. and extending this with LLMs is relatively straightforward (Gallotta et al., 2024). In parallel, LLM-based conversational agents fronted by realistic avatars with voice interfaces have a range of commercial applications, from cloning living celebrities and influencers (Bohacek and Farid, 2024; Lorenz, 2023) or deceased relatives (Lindemann, 2022; Morris and Brubaker, 2024) to offering virtual romantic companionship (Skjuve et al., 2021; Brandtzaeg et al., 2022; Pentina et al., 2023). Moving these avatars into simulated 3D environments and allowing user interactions to take place there is an obvious step.
@@
When the sort of encounter in question is with an animal, purposeful behaviour is typically discernible in the animal’s sensitivity to objects and their affordances, that is to say what they offer the agent, “for good or ill” (Gibson, 1979; Shanahan et al., 2020). The animal can be expected to interact with objects in ways that are plainly intended to fulfil its needs and achieve its goals, and to react accordingly to outside interventions, such as those of a human investigator.
The situation with virtually embodied agents of the sort envisaged is analogous. To exhibit purposeful behaviour in the context of a shared virtual world, the agent would need to do more than just talk to the user. The agent would be expected to interact with the virtual world, and with the objects it contained, in ways that were oriented towards its goals or tasks. If the agent’s behaviour were sensitive to the richness and diversity of those objects, even if they were novel, and if the agent’s ability to achieve its goals were sufficiently robust to the user’s interventions, then it would be natural to speak of its awareness of the world.
@@ 7.1 Changing attitudes
We’ll assume the model has been trained on a typically gargantuan dataset that includes sequences of actions and streams of visual input as well as the usual language corpora, and that the resulting model is embedded in a system that directs the actions of the agent’s avatar as well as what the agent says. The upshot is an embodied conversational agent with convincingly human-like behaviour.
@@ 8   Some Ethical Considerations
In the Philosophical Investigations, Wittgenstein faces down the charge of relativism from an imagined interlocutor who accuses him of saying that “human agreement decides what is true and what is false”. His reply is that humans agree in the language they use, which is “not agreement in opinions but in form of life”. [^25] PI §241. However much the language of consciousness might have to evolve to cope with the presence among us of exotic mind-like entities, our common humanity, our shared form of life, is its original home. We should strive to guide it back there whenever it strays too far.